at what temprature the kinetic energy of gas molecule is half that of its value at 27°C

Grade:11

1 Answers

Mahima Kanawat
6 years ago
KE of a gas molecule =3/2nkTSo KE is directly proportional to T (in kelvin)Therefore K /(k/2)=300/T T= 150OR you can say if KE is halved temperature will also get halved